OpenStudy (anonymous):

what is an easy way to simplify fractions (8th grade promblems)

OpenStudy (anonymous):

I think I can help you with this. What is the problem? Reduce to LCD

OpenStudy (anonymous):

just promblems in general because I have a test

OpenStudy (anonymous):

You can always break both the numerator and denominator into their composite primes. If you do that, you guarantee that it's in its simplest form if there's nothing on top that's also on bottom. If there is, you just start eliminating until there isn't. Example: 12/40 = (2*2*3)/(2*2*2*5) = 3/(2*5) = 3/10